A simple pendulum wiht length L and mass m of bob is vibrating with an amplitude of A. The maximum tension in the string is:
mg
mg(1 + (A/L)2)
mg(1+(A/2L)2)
mg(1-(3A/L)2)

Grade:

2 Answers

Arun
6 years ago
Dear student
 
Tmax = mg + mv²/l
At mean position
V = a * w and w = sqrt(g/l)
v = a sqrt (g/l)
Hence
Tmax = mg ( 1 + a²/l²)
Hence option 2 is correct.
 
Regards
Arun (askIITians forum expert)
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Tmax = mg + mv²/l
At mean position
V = a * w and w = sqrt(g/l)
v = a sqrt (g/l)
Hence
Tmax = mg ( 1 + a²/l²)

Thanks and Regards